OpenStudy (anonymous):

-3X-9=-12x+81 Solve for "x"

OpenStudy (agent47):

put all xs on one side and all ys on the other

OpenStudy (agent47):

wait i mean put all xs on one side and everything else on the other: -3x+12x=9+81 9x=90 x=10

OpenStudy (anonymous):

thanks!